OpenStudy (anonymous):

Ok it says graph all solutions on number line and provide interval notation. 1/2y + 5/4 >equal to 1/4

OpenStudy (anonymous):

\[\frac{ 1 }{ 2}y + \frac{ 5 }{ 4 }\ge \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this one

OpenStudy (anonymous):

@lgbasallote can you help me

OpenStudy (anonymous):

@Hero can u help me

OpenStudy (phi):

subtract 5/4 from both sides of the relation then multiply both sides by 2

OpenStudy (anonymous):

thanks @phi the more I sat and stared at the problem it finally clicked

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

\[y \ge -2\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

to graph, put a filled in circle at -2, and show an arrow pointing to the right